# Distributed No-Regret Learning for Multi-Stage Systems with End-to-End Bandit Feedback

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
Next, we study the multi-stage system with only end-to-end bandit feedback, that is, an agent can only observe an outcome if it receives a job and it can only observe the outcome of its chosen action. To address the exploration-exploitation-education trilemma, we propose a simple distributed online learning algorithm called $\epsilon-$EXP3. In a nutshell, the $\epsilon-$EXP3 algorithm has two operation modes, a uniform selection mode where the agent chooses actions uniformly at random to provide equal education to agents in the next stage, and an EXP3 mode where the agent employs a variation of the EXP3 algorithm to balance the tradeoff between exploration and exploitation. By randomly alternating between these two modes, the $\epsilon-$EXP3 algorithm explicitly address all three of exploration, exploitation, and education. We theoretically prove that, when applying $\epsilon-$EXP3 on a system with $L$ stages, the regret accumulated over $T$ rounds is at most $O(T^{\frac{L}{L+1}})=o(T)$.
The utility of the $\epsilon-$EXP3 algorithm is further evaluated by simulations. Simulation results show that the regret of the $\epsilon-$EXP3 algorithm indeed scales as $O(T^{\frac{L}{L+1}})$. We also evaluate two other policies that are no-regret policies for the traditional one-stage bandit problem. Surprisingly, we show that their regrets scale as $\theta(T)$ even when the system has only two stages. The simulation results demonstrate that the education component is indeed critical in multi-stage systems.
The rest of the paper is organized as follows. Section 2 surveys existing studies on adversarial bandit problems, mobile edge computing, and multi-hop networks. Section 3 introduces our system model and problem definition. Section 4 studies the simplified case with complete one-hop feedback. Section 5 introduces and analyzes the $\epsilon-$EXP3 algorithm for systems with end-to-end bandit feedback. Section 7 presents our simulation results. Finally, Section 8 concludes the paper.
@@ 2. Related Work
No-regret bandit learning. The multi-armed bandit problem has attracted significant research interests because it elegantly captures the trade-off between exploration and exploitation. In adversarial environments, the celebrated EXP3 algorithm has been proved to achieve a regret bound of $O(T^{\frac{1}{2}})$ (Auer et al., 2002). This bound has later been shown to be tight (Audibert and Bubeck, 2009). There have been many studies on variations and improvements of the EXP3 algorithm. Some studies develop variations of the EXP3 algorithm with high-probability regret bounds under both oblivious adversaries and general adversaries (Audibert and Bubeck, 2010; Neu, 2015). Some studies generalize the EXP3 algorithm for online convex optimization with bandit feedback (Flaxman et al., 2005), or for problems with multiple plays (Uchiya et al., 2010). In addition to variations of the EXP3 algorithm, there have been studies various aspects of no-regret policies, such as establishing a large familty of no-regret policies (Abernethy et al., 2015), developing algorithms whose regret bounds depend on the variance of the best arm, the path-length of the best arm, etc (Wei and Luo, 2018). All these studies only consider systems with one agent.
Mobile edge computing. Mobile edge computing (Abbas et al., 2017; Ravindran et al., 2013; Mach and Becvar, 2017) is an emerging architecture that proposes to deploy edge servers close to mobile users. One mobile edge computing application is cloud/edge robotics where a robot offloads its computation tasks to nearby edge servers. An important challenge for cloud/edge robotics is that the performance of a job depends on both the quality of the outcome and the end-to-end delay. To enable flexible trade-off between quality and latency, Jiang et al. (Jiang et al., 2018) has proposed a controller that dynamically select the suitable neural network configuration. Wu et al. (Wu et al., 2021) has modeled the problem of adaptive configuration as an integer programming problem and proposed a heuristic for it. He et al. (He et al., 2022) has employed a reinforcement learning approach for adaptive configuration. Zhang et al. (Zhang et al., 2021) has employed Lyapunov optimization to learn the optimal configuration over time. These studies only study the decisions of edge servers and they only consider stationary systems. Chinchali et al. (Chinchali et al., 2021) has proposed using deep reinforcement learning for the offloading decisions of robots, but it does not consider the adaptive configuration of edge servers. To the best of our knowledge, no existing work has jointly optimized the offloading decision of robots and the adaptive configuration of edge servers in unknown and time-varying environments.
@@ 3. System Model
Each non-leaf node $i$ employs a distributed online policy that determines the probability of forwarding a job to a child node $j$ in round $t$, denoted by $x[i,j,t]:=Prob(f[i,t]=j)$, in the event that $i$ receives a job. We have $x[i,j,t]\geq 0$ and $\sum_{j\in\mathcal{C}_{i}}x[i,j,t]=1$. Node $i$ needs to determine the values of $x[i,j,t]$ using only information available to it up to round $t-1$.
We now characterize the performance of a distributed online policy after it determines the values of $x[i,j,t]$ and selects $f[i,t]$ accordingly. Let $y[i,t]$ be the random variable indicating the amount of cost that would be incurred if node $i$ receives a job in round $t$. By definition, we have $y[j,t]=c[j,t]$ for each leaf node $j\in\mathcal{L}$. For each non-leaf node $i$, $y[i,t]$ can be calculated recursively through $y[i,t]=y[f[i,t],t]$. Also, let $w[i,t]:=E\Big{[}y[i,t]\Big{|}\mathcal{H}_{t-1}\Big{]}$ be the conditional expected amount of cost incurred if node $i$ receives a job in round $t$, given all events up to round $t-1$, denoted by $\mathcal{H}_{t-1}$. The value of $w[i,t]$ can then be calculated recursively by $w[j,t]=c[j,t]$ for each leaf node $j$ and $w[i,t]=\sum_{j\in\mathcal{C}_{i}}x[i,j,t]w[j,t]$ for each non-leaf node $i$. The total expected cost incurred by the distributed online policy over a time horizon of $T$ rounds can then be written as $\sum_{t=1}^{T}E\Big{[}y[r,t]\Big{]}=\sum_{t=1}^{T}E\Big{[}w[r,t]\Big{]}$.
@@ 5. Policy with End-to-End Bandit Feedback
$\leq$ $\min_{j\in\mathcal{C}_{i}}\sum_{t=1}^{T}y_{\epsilon}[j,t]+\epsilon_{i}T+\frac{\log D}{\eta_{i}}+\eta_{i}T\frac{D}{v[i,t]},$
@@
We now discuss the implications of Theorem 4. We can compare the $\epsilon-$EXP3 algorithm against a centrally controlled EXP3 algorithm. In the centrally controlled EXP3 algorithm, the root can control the actions of all its descendants and effectively determine which leaf node will receive a job in each round. This is equivalent to the traditional multi-armed bandit problem with $D^{L}$ arms, and the centrally controlled EXP3 algorithm can have a regret on the order of $O(\sqrt{T})$. Thus, while the $\epsilon-$EXP3 algorithm is a no-regret policy, its regret is asymptotically worse than that of the centrally controlled EXP3 algorithm. This gap is due to the need to educate children in the $\epsilon-$EXP3 algorithm. In the next section, we will establish a regret lower bound of $\Omega(T^{\frac{L-1}{L}})$ for distributed policies that need to take education into account. Since the regret of $\epsilon-$EXP3 is $O(T^{\frac{L}{L+1}})$, $\epsilon-$EXP3 is indeed near optimal.
@@ 6. Regret Lower Bound and the Need for Education
  • A2: Node $i$ makes its forwarding decision solely based on $E\big{[}y[i_{1},t]\big{]}-E\big{[}y[i_{2},t]\big{]}$. Specifically, let $\zeta:=\Big{|}E\big{[}y[i_{1},t]\big{]}-E\big{[}y[i_{2},t]\big{]}\Big{|}$, then node $i$ will forward the job to the child with higher expected cost with probability $P_{i}(\zeta)$, and to the other child with probability $1-P_{i}(\zeta)$, where $P_{i}(\cdot)$ is an arbitrary decreasing function chosen by node $i$.
We note that A1 provides a node with much more information than is possible in multi-stage systems with bandit feedback, where a node can only obtain the cost of a child if it forwards a job to the child, and only after it makes the forwarding decision. Thus, intuitively, the regret of policies with A1 serves as a natural lower bound for the regret of policies with end-to-end bandit feedback. The purpose of A2 is to highlight that a node $i$ only knows the expected costs, but not internal variables of its children.
We also note that policies with A1 do not need to explore, since it knows the expected costs of all children in advance. Hence, policies with A1 only face an education-exploitation dilemma. The only reason that a policy may select a child with higher expected cost, via choosing $P_{i}(\eta)>0$, is to educate its children.
@@ Appendix D Proof of Lemma 3
$\leq$ $\min_{j\in\mathcal{C}_{i}}\sum_{t=1}^{T}y_{\epsilon}[j,t]+\frac{\log D}{\eta_{i}}+\eta_{i}T\frac{D}{v[i,t]}.$
@@
$\leq$ $\min_{j\in\mathcal{C}_{i}}\sum_{t=1}^{T}y_{\epsilon}[j,t]+\frac{\log D}{\eta_{i}}+\eta_{i}T\frac{D}{v[i,t]}.$
@@
$\leq$ $\min_{j\in\mathcal{C}_{i}}\sum_{t=1}^{T}y_{\epsilon}[j,t]+\epsilon_{i}T+(1-\epsilon_{i})(\frac{\log D}{\eta_{i}}+\eta_{i}T\frac{D}{v[i,t]}),$
@@
$\leq$ $\min_{j\in\mathcal{C}_{i}}\sum_{t=1}^{T}y_{\epsilon}[j,t]+\epsilon_{i}T+\frac{\log D}{\eta_{i}}+\eta_{i}T\frac{D}{v[i,t]},$

### v2

@@ Abstract.
journalyear: 2024copyright: rightsretainedconference: The Twenty-fifth International Symposium on Theory, Algorithmic Foundations, and Protocol Design for Mobile Networks and Mobile Computing; October 14–17, 2024; Athens, Greecebooktitle: The Twenty-fifth International Symposium on Theory, Algorithmic Foundations, and Protocol Design for Mobile Networks and Mobile Computing (MOBIHOC ’24), October 14–17, 2024, Athens, Greecedoi: 10.1145/3641512.3686369isbn: 979-8-4007-0521-2/24/10
@@ 1. Introduction
Next, we study the multi-stage system with only end-to-end bandit feedback, that is, an agent can only observe an outcome if it receives a job and it can only observe the outcome of its chosen action. To address the exploration-exploitation-education trilemma, we propose a simple distributed online learning algorithm called $\epsilon-$EXP3. The $\epsilon-$EXP3 algorithm has two operation modes, a uniform selection mode in which the agent chooses actions uniformly at random to provide equal education to agents in the next stage, and an EXP3 mode where the agent employs a variation of the EXP3 algorithm to balance the tradeoff between exploration and exploitation. By randomly alternating between these two modes, the $\epsilon-$EXP3 algorithm explicitly address all three of exploration, exploitation, and education. We theoretically prove that, when applying $\epsilon-$EXP3 on a system with $L$ stages, the regret accumulated over $T$ rounds is at most $O(T^{\frac{L}{L+1}})=o(T)$.
To understand the fundamental regret lower bounds and the role of education in multi-stage systems, we study a class of time-homogeneous oracle policies. These policies assume that each node can know the outcome of all actions before making a decision. Therefore, there is no need to explore and each node only faces an education-exploitation dilemma. We show that the regret of these policies is at least $\Theta(T^{\frac{L-1}{L}})$, which is only slightly better than the regret of $\epsilon-$EXP3.
The utility of the $\epsilon-$EXP3 algorithm is further evaluated by simulations. The simulation results show that the regret of the $\epsilon-$EXP3 algorithm indeed scales as $O(T^{\frac{L}{L+1}})$. We also evaluate two other policies that are no-regret policies for the traditional one-stage bandit problem. Surprisingly, we show that their regrets scale as $\theta(T)$ even when the system has only two stages. The simulation results demonstrate that the education component is indeed critical in multi-stage systems.
The rest of the paper is organized as follows. Section 2 surveys existing studies on adversarial bandit problems, mobile edge computing, and multi-hop networks. Section 3 introduces our system model and problem definition. Section 4 studies the simplified case with complete one-hop feedback. Section 5 introduces and analyzes the $\epsilon-$EXP3 algorithm for systems with end-to-end bandit feedback. Section 6 establishes a regret lower bound for time-homogeneous oracle policies. Section 7 presents our simulation results. Finally, Section 8 concludes the paper.
@@ 2. Related Work
No-regret bandit learning. The multi-armed bandit problem has attracted significant research interests because it elegantly captures the trade-off between exploration and exploitation. In adversarial environments, the celebrated EXP3 algorithm has been proved to achieve a regret bound of $O(T^{\frac{1}{2}})$ (Auer et al., 2002). This bound has later been shown to be tight (Audibert and Bubeck, 2009). There have been many studies on variations and improvements of the EXP3 algorithm (Audibert and Bubeck, 2010; Neu, 2015; Uchiya et al., 2010; Wei and Luo, 2018). All these studies only consider systems with one agent.
There have been considerable recent efforts on cooperative learning (Cesa-Bianchi et al., 2020; Chawla et al., 2020; Newton et al., 2022; Agarwal et al., 2022; Landgren et al., 2021; Madhushani et al., 2021; Martínez-Rubio et al., 2019; Liu and Zhao, 2010) where agents help each other find the optimal action. These studies assume that the reward of an agent only depends on the action of that agent. In contrast, our work allows different agents to have different sets of actions and considers that the reward in each round depends on the actions of all agents. Singla, Hassani, and Krause (Singla et al., 2018) has studied a distributed learning problem in two-stage systems. It is limited to the special case of two stages and requires the root node to have the ability to block feedback information.
Mobile edge computing. One emerging application of mobile edge computing is cloud/edge robotics where a robot offloads its computation tasks to nearby edge servers. An important challenge for cloud/edge robotics is that the performance of a job depends on both the quality of the outcome and the end-to-end delay. To enable flexible trade-off between quality and latency, Jiang et al. (Jiang et al., 2018) has proposed a controller that dynamically select the suitable neural network configuration. Wu et al. (Wu et al., 2021) has modeled the problem of adaptive configuration as an integer programming problem and proposed a heuristic for it. He et al. (He et al., 2022) has employed a reinforcement learning approach for adaptive configuration. Zhang et al. (Zhang et al., 2021) has employed Lyapunov optimization to learn the optimal configuration over time. These studies only study the decisions of edge servers and they only consider stationary systems. Chinchali et al. (Chinchali et al., 2021) has proposed using deep reinforcement learning for the offloading decisions of robots, but it does not consider the adaptive configuration of edge servers. To the best of our knowledge, no existing work has jointly optimized the offloading decision of robots and the adaptive configuration of edge servers in unknown and time-varying environments.
@@ 3. System Model
Each non-leaf node $i$ employs a distributed online policy that determines the probability of forwarding a job to a child node $j$ in round $t$, denoted by $x[i,j,t]:=Prob(f[i,t]=j)$, in the event that $i$ receives a job. We have $x[i,j,t]\geq 0$ and $\sum_{j\in\mathcal{C}_{i}}x[i,j,t]=1$. Node $i$ needs to determine the values of $x[i,j,t]$ using only the information available up to round $t-1$.
We now characterize the performance of a distributed online policy after it determines the values of $x[i,j,t]$ and selects $f[i,t]$ accordingly. Let $y[i,t]$ be the random variable indicating the amount of cost that would be incurred if node $i$ receives a job in round $t$, under the probability distribution of $f[i,t]$. By definition, we have $y[j,t]=c[j,t]$ for each leaf node $j\in\mathcal{L}$. For each non-leaf node $i$, $y[i,t]$ can be calculated recursively through $y[i,t]=y[f[i,t],t]$. Also, let $w[i,t]:=E\Big{[}y[i,t]\Big{|}\mathcal{H}_{t-1}\Big{]}$ be the conditional expected amount of cost incurred if node $i$ receives a job in round $t$, given all events up to round $t-1$, denoted by $\mathcal{H}_{t-1}$. The value of $w[i,t]$ can then be calculated recursively by $w[j,t]=c[j,t]$ for each leaf node $j$ and $w[i,t]=\sum_{j\in\mathcal{C}_{i}}x[i,j,t]w[j,t]$ for each non-leaf node $i$. The total expected cost incurred by the distributed online policy over a time horizon of $T$ rounds can then be written as $\sum_{t=1}^{T}E\Big{[}y[r,t]\Big{]}=\sum_{t=1}^{T}E\Big{[}w[r,t]\Big{]}$.
@@ 5. Policy with End-to-End Bandit Feedback
$\leq$ $\min_{j\in\mathcal{C}_{i}}\sum_{t=1}^{T}y_{\epsilon}[j,t]+\epsilon_{i}T+\frac{\log D}{\eta_{i}}+\eta_{i}{\sum_{t=1}^{T}\frac{D}{v[i,t]}},$
@@ 6. Regret Lower Bound and the Need for Education
  • A2: Node $i$ makes its forwarding decision solely based on $E\big{[}y[i_{1},t]\big{]}-E\big{[}y[i_{2},t]\big{]}$. Specifically, let $\zeta:=\Big{|}E\big{[}y[i_{1},t]\big{]}-E\big{[}y[i_{2},t]\big{]}\Big{|}$, then node $i$ will forward the job to the child with the higher expected cost with probability $P_{i}(\zeta)$, and to the other child with probability $1-P_{i}(\zeta)$, where $P_{i}(\cdot)$ is an arbitrary decreasing function chosen by node $i$.
We note that A1 provides a node with much more information than is possible in multi-stage systems with bandit feedback, where a node can only obtain the cost of a child if it forwards a job to the child, and only after it makes the forwarding decision. Thus, intuitively, the regret of policies with A1 serves as a natural lower bound for the regret of policies with end-to-end bandit feedback. The purpose of A2 is to highlight that a node $i$ only knows the expected costs, but not the internal variables of its children.
We also note that policies with A1 do not need to explore, since it knows the expected costs of all children in advance. Hence, policies with A1 only face an education-exploitation dilemma. The only reason that a policy may select a child with a higher expected cost, by choosing $P_{i}(\eta)>0$, is to educate its children.
@@
Before closing the section, we note that the lower-bound analysis in this section is limited to time-homogeneous policies. We make this assumption to explicitly prevent a parent node from using history to imply internal variables of its children. Extending our analysis to time-varying policies will be interesting future work.
@@ 8. Conclusion
9. Acknowledgement
This material is based upon work supported in part by NSF under Award Numbers ECCS-2127721 and CCF-2332800 and in part by the U.S. Army Research Laboratory and the U.S. Army Research Office under Grant Number W911NF-22-1-0151.
@@ Appendix D Proof of Lemma 3
$\leq$ $\min_{j\in\mathcal{C}_{i}}\sum_{t=1}^{T}y_{\epsilon}[j,t]+\frac{\log D}{\eta_{i}}+\eta_{i}\sum_{t=1}^{T}\frac{D}{v[i,t]}.$
@@
$\leq$ $\min_{j\in\mathcal{C}_{i}}\sum_{t=1}^{T}y_{\epsilon}[j,t]+\frac{\log D}{\eta_{i}}+\eta_{i}\sum_{t=1}^{T}\frac{D}{v[i,t]}.$
@@
$\leq$ $\min_{j\in\mathcal{C}_{i}}\sum_{t=1}^{T}y_{\epsilon}[j,t]+\epsilon_{i}T+(1-\epsilon_{i})(\frac{\log D}{\eta_{i}}+\eta_{i}\sum_{t=1}^{T}\frac{D}{v[i,t]}),$
@@
$\leq$ $\min_{j\in\mathcal{C}_{i}}\sum_{t=1}^{T}y_{\epsilon}[j,t]+\epsilon_{i}T+\frac{\log D}{\eta_{i}}+\eta_{i}\sum_{t=1}^{T}\frac{D}{v[i,t]},$